### v1

# EFEAR-4D: Ego-Velocity Filtering for Efficient and Accurate 4D radar Odometry

Abstract
Odometry is a crucial component for successfully implementing autonomous navigation, relying on sensors such as cameras, LiDARs and IMUs. However, these sensors may encounter challenges in extreme weather conditions, such as snowfall and fog. The emergence of FMCW radar technology offers the potential for robust perception in adverse conditions. As the latest generation of FWCW radars, the 4D mmWave radar provides point cloud with range, azimuth, elevation, and Doppler velocity information, despite inherent sparsity and noises in the point cloud. In this paper, we propose EFEAR-4D, an accurate, highly efficient, and learning-free method for large-scale 4D radar odometry estimation. EFEAR-4D exploits Doppler velocity information delicately for robust ego-velocity estimation, resulting in a highly accurate prior guess. EFEAR-4D maintains robustness against point-cloud sparsity and noises across diverse environments through dynamic object removal and effective region-wise feature extraction. Extensive experiments on two publicly available 4D radar datasets demonstrate state-of-the-art reliability and localization accuracy of EFEAR-4D under various conditions. Furthermore, we have collected a dataset following the same route but varying installation heights of the 4D radar, emphasizing the significant impact of radar height on point cloud quality - a crucial consideration for real-world deployments. Our algorithm and dataset will be available soon at https://github.com/CLASS-Lab/EFEAR-4D.
Index Terms:
4D mmWave radar, radar odometry, ego-velocity estimation
I Introduction
Odometry estimation is a core challenge for autonomous navigation, particularly in GPS-denied environments [1, 2]. Over the past decades, vision-based and LiDAR-based odometry has been greatly advanced by the progress in sensing technologies, achieving high performance in various real-time applications [3]. However, both LiDAR and camera odometry systems still suffer from performance degradation under adverse weather conditions [1, 4, 5]. To maintain robust all-weather localization capability, several odometry methods resort to millimeter-wave (mmWave) radars as an alternative. [1, 4, 6, 7, 8]. Compared with the LiDAR, the radar exhibits a longer wavelength and is minimally affected by adverse weather conditions such as rain, snow, and fog. The ability to detect extended ranges and provide Doppler velocity information constitutes an additional advantage, facilitating the calculation of ego-velocity [9].
Compared with its predecessor 2D and 3D mmWave radars, 4D mmWave radars (called 4D radars for shortness thereafter) can provide more information in additional dimensions, that is elevation (3D radars have it as well) and Doppler velocity, but with sparser point cloud due to more compact device size and lower power consumption. These unique pros and cons necessitate developing a tailored and robust odometry solution for 4D radars. However, two challenges that need to be addressed have hindered developing effective 4D radar odometry methods. The first challenge comes from sparsity in point cloud and exclusive radar noises such as multi-path propagation and harmonics [2, 10], which make it hard to extract geometric features such as edges and planes. The second challenge comes from the limited field of view (FOV) of 4D radars, typically around $120^{\circ}$ in the horizontal plane and $40^{\circ}$ in the vertical plane, which makes reliable data association among frames challenging.
Figure 1: Our efficient 4D radar odometry and mapping.
Figure 2: Framework Overview of EFEAR-4D.
To tackle these challenges, this paper introduces a novel method for 4D radar odometry estimation, EFEAR-4D (Ego-velocity Filtering for Efficient and Accurate 4D radar Odometry), as shown in Fig. 1. Specifically, we associate direction with Doppler velocity for each point and use a density-based clustering method DBSCAN [11] to filter out noises and remove point cloud of dynamic objects, followed by ego-velocity estimation leveraging the geometric distribution characteristics of Doppler velocity from all static objects. Considering sparsity in radar point cloud, we conduct region-wise geometric and intensity-based feature extraction for registration. And Point-to-Point(P2P) registration is employed for computing odometry. Moreover, To investigate how limited FOV of 4D radars affects the numbers of point cloud and thus the performance of odometry estimation, we design experiments with 4D radar positioned at various heights, and provide comparative analysis as a guidance for future 4D radar real-world deployments. Our main contributions are summarized as follows:
  1. 1.
    A novel 4D radar odometry framework, a combination of Doppler velocity, feature-based P2P matching, is introduced. This framework effectively addresses the challenge of sparsity in 4D radar point cloud, enabling robust and precise localization in different urban environments;
  2. 2.
    An accurate method for estimating ego velocity is presented, exploiting the proposed physical model of radial Doppler velocity. Clustering and signal pre-processing algorithms are utilized to filter out dynamic objects and noises, maintaining robust performance even in the presence of complex dynamic scenes;
  3. 3.
    The proposed odometry method CFEAR-4D is thoroughly validated across various scales, vehicles, and environments, consistently demonstrating high performance in diverse scenarios. A dataset with 4D radar installation at different heights is complied. The dataset and code are open-sourced to facilitate related research in the community.
II Related Works
We will briefly summarize related works from two aspects: traditional radar (2D or 3D) and 4D imaging radar odometry.
II-A Traditional radar odometry
Radar-based odometry methods mostly can be divided into two categories: feature-based methods and direct methods. Feature-based methods typically extract and match distinctive features for registration between consecutive scans. Cen et al. [12] recognize landmarks as features to perform data association. In radarSLAM [6], SURF descriptor is extracted for pose estimation. Adolfsson et al. [1] incorporate normal vector and other factors as weights for point-to-line (P2L) scan matching. For direct methods, [13] uses the Fourier-Mellin transform to recover all rigid transformation parameters. ICP and NDT are used for radar scan registration in [14] and [15], respectively.
II-B 4D radar odometry
The Doppler velocity information of 4D radars provides a new way to estimate ego velocity that traditional 2D radars cannot. With accurate ego-velocity estimation as a reliable motion prior, pose estimation can be enhanced significantly. The simple idea of ego-velocity estimation is to consider the Doppler velocity of each stationary point as a projection of the true ego-velocity in the radial direction between the stationary point and the center of the radar. Consequently, the least squares methods can be used to calculate the ego-velocity. However, substantial noises emerge in 4D radar velocity measurements, particularly in scenarios involving dynamic objects, where the obtained Doppler velocity represents the relative velocity between two moving entities.
Kramer et al. [16] introduce Cauchy robust norm to address this challenge. Christopher and Gert [17] utilize RANSAC method to mitigate the effects of outliers. Though it works well with a small number of points, it performs poorly in large-scale scenarios, where radar point cloud typically with thousands of points. Chen et al. [8] present a 2D ego-velocity estimation method by refining ground points, however, such a method does not work in 4D cases. Due to sparsity in 4D radar point cloud, many frames lack sufficient ground points. Zeng et.al. [18] present a signal model for radar ego-motion estimation in the presence of velocity ambiguity. However, it concentrates on the fuzziness of the radar point cloud itself rather than the impact of factors such as dynamic objects in real-world situations. Therefore, we aim at proposing a novel ego-velocity filtering method, which provides accurate ego-velocity estimates by filtering out significant noises in diver dynamic scenarios.
After an accurate ego-velocity estimation, it comes to another process of registration. For 4D radar scan registration, a natural idea is to re-use methods in traditional radar or LiDAR SLAM with appropriate adjustments. APDGICP [4], a variant from generalized ICP (GICP) [19], exhibits promising performance in certain low-speed campus environments. Nevertheless, its effectiveness deteriorates significantly in high-speed driving scenarios. Li et al. [20] utilize a sliding window to build a denser radar submap and then use NDT [21] for matching. In this letter, we present a feature-based registration method adapted from CFEAR [1] 2D radar odometry and apply it to 4D radar, achieving high performance in complex environments.
III Methodology
An overview of 4D radar Odometry is shown in Fig. 2. In this section, we detail the components including ego-velocity filtering for prior guess (III-A), region-wise geometric feature extraction (III-B), and a scan-to-map matching based on motion compensation (III-C).
III-A Ego-velocity filtering
In this section, we aim to compute an accurate ego velocity as a prior guess for subsequent components of odometry estimation. The pipeline of ego-velocity filtering consists of four steps: first Gaussian filtering to filter out noises in the raw point cloud, followed by DBSCAN clustering static and dynamic points separately, then removing dynamic objects, and finally estimating ego velocity using static points.
Doppler velocity represents the relative velocity on the radial direction between the measured point and the center of the radar. Therefore, if the measured point is a stationary point, the Doppler velocity can be mathematically expressed as:
$|\textbf{v}^{R}|=\textbf{v}^{S}\cdot\textbf{r}=|\textbf{v}^{S}|\cos\theta$ (1)
where $\textbf{v}^{R}$ and $\textbf{v}^{S}$ denote the vectors of Doppler velocity and ego-velocity, $|\textbf{v}^{R}|$ denotes the value of vector $\textbf{v}^{R}$, and $r$ is the unit vector which indicates the direction of the Doppler velocity. In the default radar body frame, the $x$ and $y$ coordinates can be expressed as below:
$V_{x}$ $=|\textbf{v}^{R}|\sin\theta=\frac{|\textbf{v}^{S}|}{2}\sin{2\theta}$ (2)
$V_{y}$ $=|\textbf{v}^{R}|\cos\theta=\frac{\cos{2\theta}+1}{2}|\textbf{v}^{S}|$ (3)
As a result, we can obtain that
$V^{2}_{x}+\Big{(}V_{y}-\frac{|\textbf{v}^{S}|}{2}\Big{)}^{2}=\Big{(}\frac{|\textbf{v}^{S}|}{2}\Big{)}^{2}$ (4)
From a geometric perspective, the above equation (4) indicates that the Doppler velocity distribution for stationary points forms a circle in the XY plane and $|\textbf{v}^{S}|$ represents the diameter through the origin, as illustrated in Fig. 3. Every point on the circle can be regarded as the projection of $\textbf{v}^{S}$, reflected from a stationary place. Thus, fitting it allows us to determine the real ego velocity. In the ground vehicle setting, given that the velocity along the Z-axis is approximately zero, we directly fit a circle in the the XY plane instead of a sphere in the 3-dimensional Euclidean space. We then employ the least squares method to compute the radius $\frac{|\textbf{v}^{S}|}{2}$. The minimization function is
$\mathop{\arg\min}\limits_{r}\sum_{i}(v^{R_{x}}_{i}-0)^{2}+(v^{R_{y}}_{i}-r)^{2}-r^{2}$ (5)
where $r=\frac{|\textbf{v}^{S}|}{2}$.
Figure 3: The implementation of ego-velocity estimation. In (d), dark green points are stationary points. light green points represent the white car toward to the camera. And others are other dynamic objects or noise.
However, due to the presence of dynamic objects and noises in radar point cloud, applying the least squares method directly to all data cannot yield correct ego velocity estimates. As illustrated in Fig. 3(d), the light green points correspond to the white car driving in the opposite direction of the ego vehicle, and the dark points correspond to the pedestrians, the black car in front and some other moving objects shown in Fig. 3(a). To differentiate between dynamic points, noise, and stationary points, we employ DBSCAN [11], a density-based clustering method that utilizes region growing. Although noise and dynamic objects may not be entirely distinguishable, stationary points will dominate, allowing for the straightforward identification of stationary points and facilitating more accurate estimation. Dynamic points are then removed according to the classification results of DBSCAN for better registration in the subsequent steps.
III-B Region-wise feature extraction
Since point cloud of 4D radar is noisy, direct registration with all the points would lead to large error, especially in the Z-axis. In this section, we discuss how to extract some features for registration. Beginning with the filtered points, our initial step involves partitioning them into voxels. And for the $i$-th voxel $V_{i}$, we utilize KD-tree [22] to search for its neighboring points, which include not only points within this voxel but also adjacent points outside the area, denoted as $P_{i}$. Subsequently, employing intensity weights, we calculate the weighted mean point for $P_{i}$ as the feature point of the voxel. The weighted sample covariance matrix is denoted as $C_{i}$.
Besides, since the intensity information of 4D radar is relatively reliable, if the maximum intensity within $P_{i}$, represented as $w_{i}$, exceeds a threshold $w_{thre}$, we select the corresponding point as a feature point, which we empirically find alleviates the problem of Z-axis drift for 4D radar to a certain extent. The final feature point for $P_{i}$ is denoted as $p_{i}$. Moreover, to improve noise immunity, we perform principal component analysis (PCA) [23] to compute normal vector $u_{i}$ as another feature of the voxel $V_{i}$. The eigenvalues and eigenvectors of covariance matrix $C$ are calculated as follows:
$C_{\alpha}u_{\alpha}=\lambda_{\alpha}u_{\alpha}$ (6)
where $\alpha=1,2,3$ and $\lambda_{1}>\lambda_{2}>\lambda_{3}$. The smallest eigenvector represents $u_{i}$. The normal vectors in one scan are visualized in Fig. 4 (c). The condition number $\kappa=\lambda_{1}/\lambda_{3}$ is also calculated as a feature. To some extent, a higher value of $\kappa$ indicates greater stability and reliability of the features within this voxel.
To the best of our knowledge, it is the first time to use intensity information to extract features in 4D radar odometry. Furthermore, choosing a feature point and normal vector to represent a voxel lightens the influence of noise for registration as noise points in 4D radar point cloud are typically unstable and with low intensity.
TABLE I: QUANTITATIVE ANALYSIS: RMSE of Absolute Trajectory Error (ATE) on MSC Dataset
Method Sequence URBAN_A0 URBAN_B0 URBAN_C0 URBAN_D0 URBAN_D1 URBAN_F0 RURAL_A2 RURAL_D2
LiDAR LEGO-LOAM [24] 3.1859 33.7488 0.8671 4.4209 2.9539 9.9085 7.9991 1.0396
Vision ORB-SLAM3 [25] 11.6498 18.0216 2.2948 49.0342 52.0378 31.2399 74.3235 1.6097
[…]
4DRadarSLAM-gicp [4] 7.7078 7.6002 6.1576 92.0331 30.5685 13.5864 42.3049 0.5769
Ours-GICP 6.5796 3.4610 4.4425 57.8950 17.1938 10.9256 39.2468 1.0908
Ours 2.4282 3.2039 0.8566 15.5033 14.3756 8.4184 17.9600 1.1431
Figure 4: The current scan(a) and corresponding keyframe (b) are projected in XY plane. red points in (a) and black points in (b) are successfully matched feature points in two scans, respectively. The real current point cloud and extracted normal vectors are shown in (c). For visualization, the size of voxels here is 10m, with features being much denser in quantity during actual operation.
III-C Scan-to-map matching
Since the 4D radar point cloud are sparse, it is impossible to estimate poses reliably by matching only two consecutive frames. Therefore, we construct a keyframe sliding window for registration. Keyframes are selected based on pose deviation exceeding a predetermined threshold in rotation or translation from the preceding keyframe. The map points of keyframe within the sliding window collectively form a sub-map utilized directly in subsequent steps.
Leveraging the ego-velocity computed in the preceding step, we derive the initial guess of pose ${}^{w}\hat{T_{k}}$, and align it into sub-map coordinates for the current frame, denoted as ${}^{s}\hat{T_{k}}$. Each voxel in the current frame searches for its corresponding voxel in the sub-map in a certain radius. For corresponding voxels, feature points are then matched and a filtering process is applied to exclude points with a normal vector dot product $v_{i}\cdot v_{j}<z_{thre}$, indicating potential false association. As a result, all matching points ($\prescript{s}{}{p},\prescript{k}{}{p}$) form a sparse set $\Omega$ as shown in Fig. 4(a) and (b).
For every tuple in $\Omega$, we conduct a P2P matching. We set different weights for every tuple to distinguish some stable matching and other unstable points. The weights come from the similarity of features selected in III-B and can be calculated as follows:
$w_{i}=w^{vec}_{i}+w^{con}_{i}+w^{num}_{i}$ (7)
where $w^{vec}_{i}$,$w^{con}_{i}$ and $w^{num}_{i}$ represents the similarity of normal vectors, condition numbers, and number of points in $\prescript{s}{}{P_{i}}$ and $\prescript{k}{}{P_{i}}$, respectively. Similar to [1], here we use Huber loss as loss function due to its robustness in handling outliers. The minimization function can be formulated as follows:
$\mathop{\arg\min}\limits_{{}^{s}T_{k}}\sum_{i\in C}w_{i}L_{\delta}(||\textbf{e}_{i}||^{2})$ (8)
where $\textbf{e}_{i}=\prescript{s}{}{p}_{i}-(^{s}T_{k}\prescript{k}{}{p}_{i})$.
IV Experiments
IV-A Quantitative Evaluation
IV-A1 Comparative evaluation with MSC dataset
We start by comparing our method to previously published odometry and SLAM methods on the MSC dataset [26]. Our comparative study encompasses several vision, LiDAR, and 4D radar SLAM baselines. Specifically, ORB-SLAM3 [25], a widely recognized visual SLAM system, and LEGO-LOAM [24], a LiDAR-based SLAM approach, are selected for comparison. For fair comparison, we use only a stereo camera for ORB-SLAM3 as all baselines employ a single sensor, focusing solely on comparing odometry. For radar, we compare with 4DRadarSLAM [4], the only open-source algorithm in 4D radar slam, including two options: GICP and APDGICP. Our evaluation primarily revolves around Absolutely Trajectory Error (ATE) and Relative Pose Error (RPE) metrics [27], with RTK serving as the ground truth. The corresponding scenes are depicted in Fig. 5, encompassing static (U_A0, U_B0), high-dynamic (U_C0, U_D0, U_F0), high-speed (U_D0, U_D1), night (U_D1), snowy (R_A2, R_D2), and loop environments (U_A0, R_A2).
[…]
(b) U_B0
(c) U_C0
(d) U_D0
(e) U_D1
(f) U_F0
[…]
(h) R_D2
Figure 5: Scenes of different MSC dataset sequences. ”*” means the sequence is a loop.
(a) RURAL_A2
(b) RURAL_D2
(c) URBAN_A0
(d) URBAN_B0
(e) URBAN_D0
(f) URBAN_F0
Figure 6: Trajectories for six sequences using different SLAM algorithms on MSC dataset.
(a) URBAN_A0
(b) URBAN_D1
(c) URBAN_F0
Figure 7: Relative error of different path lengths and speeds on some sequences of MSC dataset.
We present the odometry error in Tab. I and depict the trajectory results in Fig. 6. Unexpectedly, we observe that our errors are lower than those of ORB-SLAM3 across nearly all scenarios and are approaching the level of LiDAR, or even surpassing it in some scenarios. In static environments (U_A0, U_B0), owing to precise ego velocity estimation, our EFEAR-4D method converge to a relatively accurate pose estimate, outperforming other comparative methods. In sequence U_B0, the dense vegetation makes it challenging for LiDAR to distinguish columnar features, leading to degradation in the performance of LEGO-LOAM. The other two radar methods exhibit significant errors due to the lack of feature extraction. In high-speed environment, rapid motion causes a large distance between the point cloud of consecutive frames, resulting in fewer features available for matching. As the FOV of LiDAR is much larger than radar, LEGO-LOAM demonstrates superior stability and performance. Similar to radar, the limited FOV of camera leads to rapid performance degradation of ORB-SLAM3.
In dynamic environments, by removing dynamic objects and benefiting from the weak reflection on pedestrians or other dynamic objects, our method is minimally affected by dynamic objects, unlike other methods, particularly vision-based and LiDAR-based approaches. Furthermore, compared with GICP and APDGICP, our method extracts features and filters unstable points based on intensity, further enhancing our odometry performance. In snowy environments, particularly in R_D2, where LEGO-LOAM occasionally fails to register frames, radar-based methods outperform those utilizing other sensors, demonstrating the efficacy of radar in all weather conditions.
Fig. 7 describes the relative error of ORB-SLAM3, LEGO-LOAM, baseline radar odometry, and our radar odometry EFEAR-4D using different path lengths and speeds with different sequences, following the popular KITTI odometry evaluation protocol. Our EFEAR-4D shows a stable performance on all the datasets except in high-speed scenarios where it is slightly inferior to LEGO-LOAM. Both 4DRadarSLAM (apdgicp) and ORB-SLAM3 have a relatively large error and their performances are sensitive to speed variations. It is worth pointing out that both our EFEAR-4D and 4DRadarSLAM perform well in dynamic environments, proving that 4D radar optometry is not less sensitive to moving objects than its counterparts of vision and LiDar.
IV-A2 Comparative evaluation on NTU Dataset
TABLE II: QUANTITATIVE ANALYSIS: RMSE of Relative Translation Error and Relative Rotation Error on NTU Dataset
Sequence Method livox_horizon_loam 4DRadarSLAM [4] Ours
$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
cp 0.0630 0.0195 0.0501 0.0126 0.0509 0.0125
garden 0.0563 0.0107 0.0475 0.0111 0.0379 0.0435
loop1 1.1518 0.0508 0.4300 0.0156 0.3727 0.0152
loop2 0.2217 0.0127 0.3850 0.0174 0.3701 0.0175
nyl 0.0837 0.0129 0.0909 0.0162 0.0893 0.0166
Figure 8: The point cloud and corresponding scene in NTU dataset. The majority of points are reflected from trees and bushes along the road.
We compare our methods with 4DRadarSLAM and livox_horizon_loam on NTU dataset [28]. Compared with MSC dataset, the number of sequences on NTU dataset is fewer and the scenes are less challenging since there are no dynamic objects also with relatively low speed. As the dataset only provides Livox horizon LiDAR instead of high-end scanning LiDAR, we use livox_horizon_loam as LiDAR odometry baseline. We evaluate them with the rotation part and the translation part of RPE, respectively. The results are shown in Tab. II. In loop 1 and loop 2, which are collected by car with middle speed, the drift of our EFEAR-4D in the Z-axis is much lower than that of 4DRadarSLAM due to the max-intensity point selection. In cp, a low-speed sequence where travel around buildings, our EFEAR-4D performs well as it can extract stable features such as walls and ground. However, in Garden and Nyl, where the scenes are grassland or bush, EFEAR-4D performs slightly worse as the features extracted are not distinguishable enough.
Moreover, there exists a significant difference in the number of points between the NTU dataset and the MSC dataset, with the features of points in the former appearing more ambiguous and the number of points being much lower. This is because the 4D radar is installed on the roof of the car in the NTU dataset. Owing to the limited FOV of 4D radar, point cloud are formed with reflections of objects in high spaces, such as trees. Most radar lines depicted in Fig. 8 hit trees and reflect back to the radar, exhibiting a higher trend along the road. Therefore, the points exhibit fewer distinguishable features, since trees and bushes are inherently cluttered and lack distinguishable features.
IV-A3 Ego-velocity estimation evaluation
Figure 9: Ego velocity error comparison between REVE and our method on Dynamic 2 sequence.
TABLE III: QUANTITATIVE ANALYSIS: RMSE of SPEED ERROR ON MSC DATASET
Environment Sequence Method REVE [17] Ours
$x$ $y$ $x$ $y$
Dynamic Dynamic 1 2.2864 2.2921 0.2703 0.8907
Dynamic 2 7.0582 2.0925 0.0926 0.0993
Static Static 1 0.5264 0.6038 0.2389 0.1027
Static 2 0.2068 0.3691 0.2055 0.3785
(a) Dynamic 1
(b) Dynamic 2
(c) Static 1
(d) Static 2
Figure 10: Scenes of different dynamic and static sequences.
Figure 11: DBSCAN fails to distinguish dynamic points and static object in some particular complex scenarios.
To further evaluate the performance of our ego velocity estimation, we extract 4 sequences on the MSC dataset, where two are in a dynamic environment, and the other two are in a static environment. Scenes are illustrated in Fig. 10. We compare our method with REVE [17], an open-source algorithm using RANSAC and linear least square method to compute ego velocity. As shown in Tab. III, in dynamic environments, the performance of our method is much better than REVE. The estimation errors of ego velocity in Dynamic 2 sequence are depicted in Fig. 9. Our ego velocity estimation has not only smaller errors but also smaller jitter. This is because, compared to REVE, we propose and use the physical model relating Doppler velocity to the real velocity.
However, we also observe from experiments that DBSCAN may occasionally fail to cluster in some particularly complex scenarios, as shown in Fig. 11. In such a case, all the points are used as one cluster to fit the Doppler distribution circle, resulting in a sub-optimal ego-velocity estimation.
Figure 12: Scenes and satellite imagery collected in our dataset.
TABLE IV: QUANTITATIVE ANALYSIS: RMSE of Relative Translation and Rotation Error and Absolutely ERROR on Our Dataset
Sequence Height High Low
$t_{rel}$ $t_{abs}$ $t_{rel}$ $t_{abs}$
Sequence1 0.0529 5.1802 0.0504 11.0891
Sequence2 0.0578 17.9715 0.0536 4.0270
Sequence3 0.0673 3.1901 0.0513 3.4317
(a) Sequence 1
(b) Sequence 2
(c) Sequence 3
Figure 13: The number of points in different heights with the same route.
IV-A4 Self-collected dataset
As mentioned in the previous section, the 4D radar sensor is installed on the roof of the car in the NTU dataset, its height is relatively high (compared to MSC dataset), and the overall number of points in point cloud obtained is lower than that of MSC dataset. To understand the impact of sensor height, we conduct a quantitative study: we place the 4D radar at high and low locations (about 30 cm apart) and collect sequences along the same path by traversing multiple times. Then, we test our algorithm with the same parameters with different radar height settings.
The route and scenes are depicted in Fig. 12. All the sequences are collected in University Town, Shenzhen, China. Our mobile robot platform is equipped with an Oculii Eagle 4D radar, a stereo camera, an IMU, a Robosense LiDAR, and a ByNav GNSS navigation system. Ground truth trajectory is obtained from GPS/IMU + RTK. It is clear to see that the sensor height has great impact on the number of radar points, as shown in Fig. 13. At the same GPS location, putting the radar sensor at different heights results in various signal behaviours in the urban environment. Surprisingly, the number of radar points is inconsistent at different places, going up and down from place to place.
Sequence 1 depicts the vicinity of a teaching building at night. Given its surroundings of taller buildings, its characteristics remain stable. Thus, when the radar is positioned at a higher height, it captures a large amount of building-related information, resulting in a higher count of point cloud points. Most point cloud acquired at lower heights pertain to adjacent grass areas within this sequence. Data captured at higher heights not only contains fewer points but also fewer discernible features. Insufficient points within a specific frame lead to higher drift, particularly along the Z-axis, resulting in increased errors as illustrated in Tab IV.
In Sequence 2, the point cloud contains a minimal number of points irrespective of the radar’s height, indicating a degraded radar scene characterized by the absence of discernible features. Error rates are lower when the radar is positioned at a lower height. This is because the scene lacks distinctive features and being closer to the ground enables more effective data extraction, particularly on highways. However, in open environments at higher heights, where point cloud are reflected off trees, features are lacking. Moreover, in frames with fewer than 500 points, feature extraction becomes entirely ineffective. Lowering the radar’s height facilitates better feature extraction, resulting in more accurate results as the point cloud diminishes in size.
The scenario depicted in Sequence 3 entails driving along a roadway. Despite notable differences in the point cloud density between high and low settings, the lower setting still provides adequate number of features for our EFEAR-4D to calculate the pose. In such a case, the disparity between the two settings is not significant.
V conclusion
In this paper, an 4D radar odometry framework, EFEAR-4D, is introduced. In the ego-velocity filtering module, we employ DBSCAN clustering and utilize the least square method to fit a sphere from Doppler velocity distribution for ego-velocity estimation. Furthermore, we apply region-wise feature extraction methods to extract stable features. P2P matching is enhanced by adding a keyframe sliding window and using Huber function. Additionally, a new dataset is collected. We conduct our experiments on two public datasets and our self-collected dataset. Our method outperforms the state-of-the-art odometry methods with vision or LiDAR in different challenging environments, making 4D radar an attractive option for robust and accurate localization.
Some findings in our experiment show that 4D radar may show some weakness in terms of perception for certain scenarios. However, the shortcoming are expected to be addressed with the integration of other complementary sensors. In the future, we will explore sensor fusion with 4D radar and cameras to provide a more comprehensive odometry solution.